OpenStudy (amistre64):

What is the equation of the tangent plane to the function: 2x^2 +y^2 +2z^2 +2xz =7. At the point (-1,1,2) ?

OpenStudy (amistre64):

it begins with the gradient....

OpenStudy (anonymous):

yeah thatl giv u d direction ratios of the tangent plane

OpenStudy (anonymous):

4x+2z, 2y, 4z+2x thats 0, 2, 6

OpenStudy (amistre64):

goody goody goody :)

OpenStudy (anonymous):

but then how do u get the direction ratios of the normal?

OpenStudy (amistre64):

the normal is <0,2,6>; so use the point to calibrate the equation

OpenStudy (anonymous):

thats the ratios parallel to the plane

OpenStudy (amistre64):

0(x+1)+2(y-1)+6(z-2)=0

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

those coefficients have to be in the direction of the normal, and not the tangent

OpenStudy (amistre64):

can you explain?

OpenStudy (amistre64):

n=<a,b,c>; P=(x0,y0,z0) equation of the plane: a(x-x0)+b(y-y0)+c(z-z0)=0 is the way the textbook tells me it

OpenStudy (anonymous):

ur using the standard eqn of a plane passing through a given pt a,b,c A(x-a) + B(x-b) +C(x-c)=0 where A,B,C are direction ratios of the NORMAL

OpenStudy (amistre64):

x-x0; y-y0 and z-z0 are the components of the vector on the plane

OpenStudy (anonymous):

u get the cross product of <0,2,6> and <-1, 1,2>

OpenStudy (anonymous):

thatl give u a vector normal to the plane say <A,B,C>

OpenStudy (anonymous):

then use the standard equation

OpenStudy (amistre64):

-1,1,2 is a point on the plane right? not a vector parallel to it tho.

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

the vector that is parallel to the plane is (x-x0,y-y0,z-z0) where P(x0,y0,z0) is on the plane right?

OpenStudy (anonymous):

-2(x+1) +6(y-1) +2 (z-2) = 0

OpenStudy (amistre64):

-2,6,2 is not the normal to the tangent plane tho. if anything it too is parallel to the plane

OpenStudy (anonymous):

arent you suppose to find Fx(X,Y) and Fy(X,Y) use the cross product to find the N?

OpenStudy (anonymous):

i told u get the vector product of the direction of the tangent and the point

OpenStudy (anonymous):

weve already found the gradient, i suggest we cross that with the point to get a vector normal to the plane

OpenStudy (amistre64):

the normal was found by the gradient; isnt the gradient and the normal parallel vectors?

OpenStudy (anonymous):

no the gradient is parallel to the plane

OpenStudy (anonymous):

If you found the N and have a point whats the problem?

OpenStudy (amistre64):

hmmm.... whats the gradient of 0 = 2x -y +4

OpenStudy (amistre64):

just trying to understand concepts :)

OpenStudy (amistre64):

<2,-1,0> right?

OpenStudy (amistre64):

is that perp or // to y=2x+4 ; z=0?

OpenStudy (anonymous):

||

OpenStudy (amistre64):

so the vector <2,-1>; whose slope is -1/2 is // with the plne y=2x+2; z=0?

OpenStudy (amistre64):

(-1/2)*2=-1.....

OpenStudy (anonymous):

u cant define a gradient vector for a line

OpenStudy (amistre64):

i didnt; i defined a plane; hence the z=0

OpenStudy (anonymous):

fine

OpenStudy (anonymous):

bt m saying i think wt i did ws right

OpenStudy (anonymous):

Do you guys know about green theorem?

OpenStudy (amistre64):

i can tell :)

OpenStudy (anonymous):

ive heard f it..nt studied it..

OpenStudy (anonymous):

hey sorry..i confirmed..the gradient vector is normal to the surface..rly sorry

OpenStudy (amistre64):

if we dot the gradient to a vector parallel to the plane y=2x+4; z=0, we will know if its perp or //

OpenStudy (anonymous):

no i confirmed

OpenStudy (amistre64):

g=<2,-1,0> v=<1, 2,0> -------- 2-2+0=0

OpenStudy (anonymous):

yes m tellin u its normal to the surface

OpenStudy (anonymous):

wt u were doing is right...i feel u alrdy know all this nd ur just testing us out

OpenStudy (amistre64):

;) gotta keep entertained ya know lol

OpenStudy (anonymous):

mean guy

OpenStudy (amistre64):

id give you another medal if i could ;)

OpenStudy (anonymous):

wts stopping u??

OpenStudy (amistre64):

the only option available to me is the undo lol